# Finding Flawed Fictions: Evaluating Complex Reasoning in Language Models via Plot Hole Detection

### v2

@@ 1 Introduction
Narratives form a fundamental mode of human cognition and meaning-making, acting as a primary way people organize, experience, and construct reality (Bruner, 1991). When we engage with stories, we typically go beyond a literal understanding of what happened, instead performing complex and nuanced reasoning that involves mental representation of a story’s world and its characters (Gerrig, 1993; Mar & Oatley, 2008; Zunshine, 2006; Kidd & Castano, 2013). Ultimately, narrative understanding is a reflection of broader human cognitive capacities for language comprehension and reasoning (Kintsch, 1998).
@@
In this work, we propose to quantify narrative understanding in LLMs as a novel test bed of general language understanding and reasoning abilities. While different language understanding benchmarks are widespread in existing literature(Wang et al., 2018; 2019; Zellers et al., 2019; Hendrycks et al., 2020; Jaradeh et al., 2023), they often fail to capture the full spectrum of abilities present in narrative understanding. For example, the popular MMLU benchmark (Hendrycks et al., 2020) evaluates advanced multi-hop knowledge, but lacks assessment of pragmatics and implicit social dynamics inherent in narratives. Existing datasets studying such capabilities (Mostafazadeh et al., 2016; Sap et al., 2019; Sprague et al., 2024; Kim et al., 2023), on the other hand, are not suited for benchmarking LLMs at scale, as they focus on very short or fully synthetic stories that lack core elements of narrative structure. As a consequence, it remains difficult to holistically assess overall progress in language understanding and reasoning, despite recent advances in improving LLM reasoning capabilities through advanced prompting (Wei et al., 2022; Yao et al., 2024; Wang et al., 2023) or inference time scaling (Lambert et al., 2024; Guo et al., 2025).
@@
How do we quantify such “deeper narrative understanding”? We propose a novel task of plot hole detection as a proxy to assess deep narrative understanding and reasoning in LLMs. Plot holes are inconsistencies in a story that go against the logic flow established by the story plot (Ryan, 2009), with significant discourse dedicated to both locating [^1] This is especially true in the context of films, with dedicated subreddits like r/plotholes and r/MovieMistakes, or a Goofs section dedicated to each film page on IMDB. and preventing them during screen writing (McKee, 1997; MasterClass, 2021). Plot hole detection requires nuanced reasoning about the implications of established facts and elements, how they interplay, and their plausibility. Specifically, robust state tracking is needed to follow entities and rules established by the story over a long context; commonsense and pragmatic reasoning are needed for interpreting implicit world knowledge and beliefs; and theory of mind is required for reasoning over beliefs, motivations, and desires of characters. Beyond acting as a test bed for complex reasoning, models that can accurately assess plot holes in stories can be useful to improve consistency in writing, be it human- or machine-generated.
@@ 2 Defining Plot Holes: Continuity Errors
Plot holes are commonly categorized into multiple categories (Shattuck, 2024) including: continuity errors (contradictions of established facts), out of character behavior (actions inconsistent with established motivations), factual errors (historical anachronisms or real-world inaccuracies), impossible events (violations of science or logic), and unresolved storylines (incomplete plot threads). See Table LABEL:table:plotholeexamples in Appendix for examples. We focus on continuity errors as they encompass the most general form of plot hole: both out of character behavior and impossible events can be framed as breaks in continuity, as they contradict established character traits or story settings. While Ryan (2009) distinguishes between harmless plot holes (serving symbolic functions rather than causal functions) and truly unbridgeable ones (affecting plot integrity), our approach treats both types as under the same umbrella.
Formally, consider a fictional story $f$ containing a set of propositions =$\mathcal{F}=\{\phi_{1},\ldots,\phi_{n}\}$ that are true in the fictional world of $f$ (e.g., “Sherlock Holmes lived on Baker Street” is a statement that is true in the fictional world of Sherlock Holmes). We make use of the possible worlds theory from Lewis (1978), defining the notation $\text{isTrue}(f,\phi)$ to denote that the proposition $\phi$ is true in the fictional world of $f$ and define the shorthand $\text{isTrue}(f,\mathcal{F})~{}\!:=~{}\!\text{isTrue}(f,\phi_{1})~{}\!\wedge~{}\!\cdots~{}\!\wedge~{}\!\text{isTrue}(f,\phi_{n})$. We can then define a continuity error:
@@ 3 Automatically Generating Plot Holes in Stories
1. Three Act Structure Extraction. We start by dividing the story into the traditional three act structure Aristotle (1902), consisting of Act One ($A_{1}$), where the main characters and setting of the story are introduced, Act Two ($A_{2}$), where the main conflict is developed, and Act Three ($A_{3}$), which builds to the climax and resolves the main conflict. This division aids to control where the original proposition is established in the story and when it gets contradicted in the later parts of our pipeline. We perform the three-act extraction of an original story $f$ through LLM prompting, and denote it $\{A_{1},A_{2},A_{3}\}\leftarrow\texttt{ThreeActExtract}(f)$. Note that $f$ is the concatenation $f=A_{1}\cdot A_{2}\cdot A_{3}$ of the resulting three acts $\{A_{1},A_{2},A_{3}\}$.
@@
3. Counterfactual Story Generation. We rewrite the story while negating an original proposition $\phi$ with LLM prompting (Qin et al., 2019), $A_{1}^{\neg\phi}\!\cdot\!A_{2}^{\neg\phi}\!\cdot\!A_{3}^{\neg\phi}\!\leftarrow\!\texttt{Counterfact}(\phi,A_{1},A_{2},A_{3})$. Note that negating $\phi$ does not just negate that single statement in the story, but may also lead to modifying other existing propositions to maintain coherence and plausibility (e.g., when changing a character’s nationality, their name might need to be changed).
@@
5. Filtering. As a final step, we ensure that the patched story results in an inherent story inconsistency. This includes removing obvious LLM prompting issues, such as cases where $A_{2}=A_{2}^{\neg\phi}$ or $A_{3}=A_{3}^{\neg\phi}$, or preemptively removing cases where there are too many changes ($>5$) in the counterfactual, since an increasing number of LLM edits increases the probability of making counterfactual reasoning errors. We additionally run an extremely aided version of the task as a quality filter: we prompt an LLM with $f^{\text{patch}}$, specifying the modified lines in $A_{2}^{\neg\phi}$ and $A_{3}^{\neg\phi}$ and use the LLM as a judge of whether these lines introduce a continuity error. This much simpler problem [^4] This is a much simpler problem because the model only needs to check the lines marked for a contradiction, as opposed to all the possible combinations of them. aids us in eliminating cases with errors during Step 3, where the newly introduced propositions might still be consistent with the original fact $\phi$. To improve reliability of filtering, we use self-consistency (Wang et al., 2023), only retaining the cases where the model predicts a continuity error in at least 4 out of the 5 completions. At the filtering step we also prompt the model to provide an explanation if it predicts that the modified lines introduce a continuity error, which is shown later to humans to verify if the stories actually have a continuity error.
We use GPT-4o for all steps, except for counterfactual story generation where we qualitatively found GPT-4-turbo to perform significantly better. All the prompts used for our pipeline are provided in Appendix § A.10.1. While four out of five steps in our pipeline make use of LLMs, we do not claim that LLMs to be perfect at these tasks. Step 3, which requires counterfactual reasoning can in particular be difficult for LLMs with evidence in prior work (Huang et al., 2024). Hence, we follow our automatic generation process with human verification to curate a high quality benchmark.
@@ 4 FlawedFictions: Tasks, Metrics, and Dataset Statistics
FlawedFictionsLong. Our preliminary experiments showed LLMs struggle with assessing plot holes as story length increased (see §A.5.2 in Appendix). Consequently, we curated an extension of FlawedFictionsFlawedFictionsLong – consisting of stories 1,200-4,000 words long: we selected stories from FairyTaleQA (Xu et al., 2022) meeting this length criterion and processed them through FlawedFictionsMaker to generate positive examples. Due to increased cognitive load and annotation costs, only one-third of these longer stories were annotated by Prolific users, with the remainder annotated by this paper’s lead author. Post-verification, we selected 97 stories as positive examples and 103 original stories as negative examples, totaling 200 examples in FlawedFictionsLong. Unlike FlawedFictions, FlawedFictionsLong consists entirely of fairy tales and has an average length of 2703 words per story.
@@ 5 How Well do Frontier LLMs Perform on FlawedFictions?
Experimental Setup. We evaluate different proprietary LLMs from OpenAI and Anthropic as well as open weights models Llama-3 (Van Der Maaten et al., 2024), Deepseek-R1 Distilled (Guo et al., 2025), and Qwen-2.5 (Yang et al., 2024) series, which represent the most recent iterations available at the time of publication. For o1 and o3-mini, we experiment with the three values of reasoning efforts parameter provided in the API – low, medium, and high, which controls the amount of intermediate reasoning tokens generated before the final completion. Similarly, Anthropic API provides extended thinking mode for Claude 3.7 Sonnet model, which uses intermediate tokens to “think” before answering. We also consider another inference time scaling strategy, where we augment the plot hole detection model i.e. generator with a verifier model (Cobbe et al., 2021) that validates the legitimacy of the plot hole detected by the generator. Our verifier is a Claude 3.5 Sonnet model prompted to perform the verification task. For more details on the experimental setup, prompts that we use, and other prompting methods that we evaluate such as few-shot and chain-of-thought (CoT), refer to Appendix §A.4.
Baselines. To highlight the contextual nature of our problem, we use an entailment model that examines all ordered sentence pairs in a story to detect contradictions. If no contradictory pairs are found, the baseline predicts the story lacks continuity errors; otherwise, the pair with highest contradiction confidence determines the error location. We employ DeBERTa-v3-large (He et al., 2021) fine-tuned on MNLI (Williams et al., 2018) (achieving 91% on MNLI dev) as our entailment model. We also consider a random baseline and a baseline that always predicts No continuity error found, with the latter achieving 50% on CEEval-Full due to our balanced dataset.
Benchmarking Human Performance. To establish a meaningful baseline against which to compare performance of various LLMs on FlawedFictions, we estimated human performance by recruiting 9 undergraduate English majors who evaluated 50 samples from FlawedFictions with three responses per sample. Further details about the study are provided in Appendix §A.2. It is important to recognize that this task is non-trivial for humans as it requires a high amount of cognitive load due to the limited working memory, which has been shown to affect reading comprehension abilities in adults and children (Barreyro et al., 2025; Cain et al., 2004).
@@ 6 Measuring Logical Consistency in LLM Generated Narratives
A study by Mirowski et al. (2023) examining LLMs as screenplay co-writers identified that LLM-generated narratives exhibited issues with maintaining consistency in plot’s logic or characters’ behaviors. While these observations were made based on participants’ interviews, we propose a quantitative evaluation framework for the phenomenon. Our setup consists of generating short stories using LLMs, which are subsequently evaluated for the existence of plot holes using our best model on FlawedFictions i.e. Claude 3.5 Sonnet with Verifier. We define continuity error detection rate as the fraction of the generated stories for which the detection model identifies a continuity error.
@@ 7 Related Work
Narrative Understanding and Reasoning Tasks. Narrative understanding tasks can be categorized as descriptive or interpretive. Descriptive tasks, which involve understanding explicitly stated plot elements, include question answering benchmarks (NarrativeQA (Kočiský et al., 2018), FairyTaleQA (Xu et al., 2022), and BookQA (Angelidis et al., 2019)), narrative summarization (Ouyang et al., 2017; Papalampidi et al., 2020; Kryscinski et al., 2022), and claim verification (Karpinska et al., 2024). Interpretive tasks require forming mental representation of story’s worlds and utilzing those to infer their logical implications, such as selecting correct endings (Mostafazadeh et al., 2016), assessing causality (Roemmele et al., 2011), or generating counterfactuals (Qin et al., 2019). However, unlike FlawedFictions, these datasets focus on very short stories that are roughly 4 to 5 sentences long. While, MuSR (Sprague et al., 2024) introduced multi-step reasoning over narratives involving tasks like solving murder mysteries, it uses synthetic stories with specific templates, whereas FlawedFictions comprises edited versions of human-written stories with diverse narrative structures.
@@
Evaluating Quality of LLM Generated Stories. Studies show GPT-3-generated stories score highly on fluency and coherence compared to specifically tuned models and competitively with humans (Xie et al., 2023). However, human-written stories have been shown to exhibit more diverse narrative structures than the largely homogeneous LLM-generated stories (Tian et al., 2024). While GPT-4 stories surpass human-written ones on the Psychological Depth Scale (Harel-Canada et al., 2024), which quantifies the emotion, empathy, and engagement in stories, they score lower on the Creativity Index (Lu et al., 2025), which measures linguistic creativity by searching for verbatim matches against web documents. None of these measure the logical and motivational consistency of narratives and there is evidence (Mirowski et al., 2023) that LLM authored stories can lack plot and character consistency.
@@
Plot Holes and Impossible Worlds. Plot holes are inadvertent inconsistencies in a story’s logical and motivational texture (Ryan, 2009). Lewis (1978) defines such stories where the plot contradicts itself as impossible fictions, citing the example of contradicting locations of Watson’s old war wound in Sherlock Holmes. Lewis (1978) proposes resolutions of truth in such fictions by considering revisions that remain close to the original. Badura & Berto (2019) extends this theory with “impossible worlds” that can contain logical contradictions without rendering everything vacuously true to make sense of stories that deliberately defy logic (Priest, 1997). Plot holes have also been discussed in mathematics education contexts (Miežys, 2023).
@@
Automatic Detection of Plot Holes. Davids (2022) introduced a symbolic approach using epistemic logic to identify plot holes, though the approach requires structured story events and is not flexible to operate on any story. Chadalapaka et al. (2023) generate synthetic data for plot hole detection by negating a randomly sampled statement in the story. However, this approach may not consistently generate plot holes, and to the best of our knowledge the authors do not perform human verification for their generated data.

### v3

@@ 1 Introduction
Narratives form a fundamental mode of human cognition and meaning-making, acting as a primary way people organize, experience, and construct reality (Bruner1991-BRUTNC). When we engage with stories, we typically go beyond a literal understanding of what happened, instead performing complex and nuanced reasoning that involves mental representation of a story’s world and its characters (gerrig1993experiencing; Mar2008TheFO; zunshine2006why; kidd2013reading). Ultimately, narrative understanding is a reflection of broader human cognitive capacities for language comprehension and reasoning (kintsch1998comprehension).
@@
In this work, we propose to quantify narrative understanding in LLMs as a novel test bed of general language understanding and reasoning abilities. While different language understanding benchmarks are widespread in existing literature(wang-etal-2018-glue; wang2019superglue; zellers2019hellaswag; hendrycks2020measuring; jaradeh2023sciqa), they often fail to capture the full spectrum of abilities present in narrative understanding. For example, the popular MMLU benchmark (hendrycks2020measuring) evaluates advanced multi-hop knowledge, but lacks assessment of pragmatics and implicit social dynamics inherent in narratives. Existing datasets studying such capabilities (mostafazadeh-etal-2016-corpus; sap-etal-2019-social; sprague2024musr; kim-etal-2023-fantom), on the other hand, are not suited for benchmarking LLMs at scale, as they focus on very short or fully synthetic stories that lack core elements of narrative structure. As a consequence, it remains difficult to holistically assess overall progress in language understanding and reasoning, despite recent advances in improving LLM reasoning capabilities through advanced prompting (wei2022chain; yao2024tree; wang2023selfconsistency) or inference time scaling (lambert2024t; guo2025deepseek).
@@
How do we quantify such “deeper narrative understanding”? We propose a novel task of plot hole detection as a proxy to assess deep narrative understanding and reasoning in LLMs. Plot holes are inconsistencies in a story that go against the logic flow established by the story plot (ryan2009cheap), with significant discourse dedicated to both locating [^1] This is especially true in the context of films, with dedicated subreddits like r/plotholes and r/MovieMistakes, or a Goofs section dedicated to each film page on IMDB. and preventing them during screen writing (McKee1997; masterclass_plotholes). Plot hole detection requires nuanced reasoning about the implications of established facts and elements, how they interplay, and their plausibility. Specifically, robust state tracking is needed to follow entities and rules established by the story over a long context; commonsense and pragmatic reasoning are needed for interpreting implicit world knowledge and beliefs; and theory of mind is required for reasoning over beliefs, motivations, and desires of characters. Beyond acting as a test bed for complex reasoning, models that can accurately assess plot holes in stories can be useful to improve consistency in writing, be it human- or machine-generated.
@@ 2 Defining Plot Holes: Continuity Errors
Plot holes are commonly categorized into multiple categories (shattuck-2024-six) including: continuity errors (contradictions of established facts), out of character behavior (actions inconsistent with established motivations), factual errors (historical anachronisms or real-world inaccuracies), impossible events (violations of science or logic), and unresolved storylines (incomplete plot threads). See Table LABEL:table:plotholeexamples in Appendix for examples. We focus on continuity errors as they encompass the most general form of plot hole: both out of character behavior and impossible events can be framed as breaks in continuity, as they contradict established character traits or story settings. While ryan2009cheap distinguishes between harmless plot holes (serving symbolic functions rather than causal functions) and truly unbridgeable ones (affecting plot integrity), our approach treats both types as under the same umbrella.
Formally, consider a fictional story $f$ containing a set of propositions $\mathcal{F}=\{\phi_{1},\ldots,\phi_{n}\}$ that are true in the fictional world of $f$ (e.g., “Sherlock Holmes lived on Baker Street” is a statement that is true in the fictional world of Sherlock Holmes). We make use of the possible worlds theory from lewis1978truth, defining the notation $\text{isTrue}(f,\phi)$ to denote that the proposition $\phi$ is true in the fictional world of $f$ and define the shorthand $\text{isTrue}(f,\mathcal{F})~\!:=~\!\text{isTrue}(f,\phi_{1})~\!\wedge~\!\cdots~\!\wedge~\!\text{isTrue}(f,\phi_{n})$. We can then define a continuity error:
@@ 3 Automatically Generating Plot Holes in Stories
1. Three Act Structure Extraction. We start by dividing the story into the traditional three act structure aristotle1902poetics, consisting of Act One ($A_{1}$), where the main characters and setting of the story are introduced, Act Two ($A_{2}$), where the main conflict is developed, and Act Three ($A_{3}$), which builds to the climax and resolves the main conflict. This division aids to control where the original proposition is established in the story and when it gets contradicted in the later parts of our pipeline. We perform the three-act extraction of an original story $f$ through LLM prompting, and denote it $\{A_{1},A_{2},A_{3}\}\leftarrow\texttt{ThreeActExtract}(f)$. Note that $f$ is the concatenation $f=A_{1}\cdot A_{2}\cdot A_{3}$ of the resulting three acts $\{A_{1},A_{2},A_{3}\}$.
@@
3. Counterfactual Story Generation. We rewrite the story while negating an original proposition $\phi$ with LLM prompting (Qin2019CounterfactualSR), $A_{1}^{\neg\phi}\!\cdot\!A_{2}^{\neg\phi}\!\cdot\!A_{3}^{\neg\phi}\!\leftarrow\!\texttt{Counterfact}(\phi,A_{1},A_{2},A_{3})$. Note that negating $\phi$ does not just negate that single statement in the story, but may also lead to modifying other existing propositions to maintain coherence and plausibility (e.g., when changing a character’s nationality, their name might need to be changed).
@@
5. Filtering. As a final step, we ensure that the patched story results in an inherent story inconsistency. This includes removing obvious LLM prompting issues, such as cases where $A_{2}=A_{2}^{\neg\phi}$ or $A_{3}=A_{3}^{\neg\phi}$, or preemptively removing cases where there are too many changes (>5>5) in the counterfactual, since an increasing number of LLM edits increases the probability of making counterfactual reasoning errors. We additionally run an extremely aided version of the task as a quality filter: we prompt an LLM with $f^{\text{patch}}$, specifying the modified lines in $A_{2}^{\neg\phi}$ and $A_{3}^{\neg\phi}$ and use the LLM as a judge of whether these lines introduce a continuity error. This much simpler problem [^4] This is a much simpler problem because the model only needs to check the lines marked for a contradiction, as opposed to all the possible combinations of them. aids us in eliminating cases with errors during Step 3, where the newly introduced propositions might still be consistent with the original fact $\phi$. To improve reliability of filtering, we use self-consistency (wang2023selfconsistency), only retaining the cases where the model predicts a continuity error in at least 4 out of the 5 completions. At the filtering step we also prompt the model to provide an explanation if it predicts that the modified lines introduce a continuity error, which is shown later to humans to verify if the stories actually have a continuity error.
We use GPT-4o for all steps, except for counterfactual story generation where we qualitatively found GPT-4-turbo to perform significantly better. All the prompts used for our pipeline are provided in Appendix § A.10.1. While four out of five steps in our pipeline make use of LLMs, we do not claim that LLMs to be perfect at these tasks. Step 3, which requires counterfactual reasoning can in particular be difficult for LLMs with evidence in prior work (huang-etal-2024-clomo). Hence, we follow our automatic generation process with human verification to curate a high quality benchmark.
@@ 4 FlawedFictions: Tasks, Metrics, and Dataset Statistics
FlawedFictionsLong. Our preliminary experiments showed LLMs struggle with assessing plot holes as story length increased (see §A.5.2 in Appendix). Consequently, we curated an extension of FlawedFictionsFlawedFictionsLong – consisting of stories 1,200-4,000 words long: we selected stories from FairyTaleQA (xu-etal-2022-fantastic) meeting this length criterion and processed them through FlawedFictionsMaker to generate positive examples. Due to increased cognitive load and annotation costs, only one-third of these longer stories were annotated by Prolific users, with the remainder annotated by this paper’s lead author. Post-verification, we selected 97 stories as positive examples and 103 original stories as negative examples, totaling 200 examples in FlawedFictionsLong. Unlike FlawedFictions, FlawedFictionsLong consists entirely of fairy tales and has an average length of 2703 words per story.
@@ 5 How Well do Frontier LLMs Perform on FlawedFictions?
Experimental Setup. We evaluate different proprietary LLMs from OpenAI and Anthropic as well as open weights models Llama-3 (vandermaaten2024llama3), Deepseek-R1 Distilled (guo2025deepseek), and Qwen-2.5 (qwen2.5) series, which represent the most recent iterations available at the time of publication. For o1 and o3-mini, we experiment with the three values of reasoning efforts parameter provided in the API – low, medium, and high, which controls the amount of intermediate reasoning tokens generated before the final completion. Similarly, Anthropic API provides extended thinking mode for Claude 3.7 Sonnet model, which uses intermediate tokens to “think” before answering. We also consider another inference time scaling strategy, where we augment the plot hole detection model i.e. generator with a verifier model (Cobbe2021TrainingVT) that validates the legitimacy of the plot hole detected by the generator. Our verifier is a Claude 3.5 Sonnet model prompted to perform the verification task. For more details on the experimental setup, prompts that we use, and other prompting methods that we evaluate such as few-shot and chain-of-thought (CoT), refer to Appendix §A.4.
Baselines. To highlight the contextual nature of our problem, we use an entailment model that examines all ordered sentence pairs in a story to detect contradictions. If no contradictory pairs are found, the baseline predicts the story lacks continuity errors; otherwise, the pair with highest contradiction confidence determines the error location. We employ DeBERTa-v3-large (he2021deberta) fine-tuned on MNLI (williams-etal-2018-broad) (achieving 91% on MNLI dev) as our entailment model. We also consider a random baseline and a baseline that always predicts No continuity error found, with the latter achieving 50% on CEEval-Full due to our balanced dataset.
Benchmarking Human Performance. To establish a meaningful baseline against which to compare performance of various LLMs on FlawedFictions, we estimated human performance by recruiting 9 undergraduate English majors who evaluated 50 samples from FlawedFictions with three responses per sample. Further details about the study are provided in Appendix §A.2. It is important to recognize that this task is non-trivial for humans as it requires a high amount of cognitive load due to the limited working memory, which has been shown to affect reading comprehension abilities in adults and children (Barreyro2025; Cain2004).
@@ 6 Measuring Logical Consistency in LLM Generated Narratives
A study by mirowski2023co examining LLMs as screenplay co-writers identified that LLM-generated narratives exhibited issues with maintaining consistency in plot’s logic or characters’ behaviors. While these observations were made based on participants’ interviews, we propose a quantitative evaluation framework for the phenomenon. Our setup consists of generating short stories using LLMs, which are subsequently evaluated for the existence of plot holes using our best model on FlawedFictions i.e. Claude 3.5 Sonnet with Verifier. We define continuity error detection rate as the fraction of the generated stories for which the detection model identifies a continuity error.
@@ 7 Related Work
Narrative Understanding and Reasoning Tasks. Narrative understanding tasks can be categorized as descriptive or interpretive. Descriptive tasks, which involve understanding explicitly stated plot elements, include question answering benchmarks (NarrativeQA (kocisky-etal-2018-narrativeqa), FairyTaleQA (xu-etal-2022-fantastic), and BookQA (angelidis-etal-2019-book)), narrative summarization (ouyang-etal-2017-crowd; papalampidi-etal-2020-screenplay; kryscinski-etal-2022-booksum), and claim verification (karpinska-etal-2024-one). Interpretive tasks require forming mental representation of story’s worlds and utilzing those to infer their logical implications, such as selecting correct endings (mostafazadeh-etal-2016-corpus), assessing causality (roemmele_choice_2011), or generating counterfactuals (Qin2019CounterfactualSR). However, unlike FlawedFictions, these datasets focus on very short stories that are roughly 4 to 5 sentences long. While, MuSR (sprague2024musr) introduced multi-step reasoning over narratives involving tasks like solving murder mysteries, it uses synthetic stories with specific templates, whereas FlawedFictions comprises edited versions of human-written stories with diverse narrative structures. For a review about narrative theory for narrative understanding, please refer to piper-etal-2021-narrative.
@@
Evaluating Quality of LLM Generated Stories. Studies show GPT-3-generated stories score highly on fluency and coherence compared to specifically tuned models and competitively with humans (xie-etal-2023-next). However, human-written stories have been shown to exhibit more diverse narrative structures than the largely homogeneous LLM-generated stories (tian-etal-2024-large-language). While GPT-4 stories surpass human-written ones on the Psychological Depth Scale (harel-canada-etal-2024-measuring), which quantifies the emotion, empathy, and engagement in stories, they score lower on the Creativity Index (lu2025ai), which measures linguistic creativity by searching for verbatim matches against web documents. None of these measure the logical and motivational consistency of narratives and there is evidence (mirowski2023co) that LLM authored stories can lack plot and character consistency.
@@
Plot Holes and Impossible Worlds. Plot holes are inadvertent inconsistencies in a story’s logical and motivational texture (ryan2009cheap). lewis1978truth defines such stories where the plot contradicts itself as impossible fictions, citing the example of contradicting locations of Watson’s old war wound in Sherlock Holmes. lewis1978truth proposes resolutions of truth in such fictions by considering revisions that remain close to the original. Badura02012019 extends this theory with “impossible worlds” that can contain logical contradictions without rendering everything vacuously true to make sense of stories that deliberately defy logic (priest1997sylvan). Plot holes have also been discussed in mathematics education contexts (Miezys2023).
@@
Automatic Detection of Plot Holes. essay91967 introduced a symbolic approach using epistemic logic to identify plot holes, though the approach requires structured story events and is not flexible to operate on any story. chadalapaka2023low generate synthetic data for plot hole detection by negating a randomly sampled statement in the story. However, this approach may not consistently generate plot holes, and to the best of our knowledge the authors do not perform human verification for their generated data.